Farrelly enterprises has fixed operating cost of $500,000, variable operating cost per unit of $20 and a selling price of $40. The firm's capital structure consists of $600,000 loan at 10% interest, 10,000 shares of preferred stock paying an annual dividend of $3 per share and $50,000 shares of common stock outstanding. Farrelly has a 34% tax rate. Calculator Farrelly operating breakeven point in units and in dollar sales
1
Expert's answer
2015-02-25T09:30:55-0500
FC = $500,000, VC per unit = $20, P = $40, $600,000 loan at 10% interest, 10,000 shares of preferred stock (dividends $3 per share),
$50,000 shares of common stock outstanding, T = 34%.
Breakeven Point in Units = FC/(P - VC) = 500,000/(40 - 20) = 25,000 units
Breakeven Point in dollar sales = 25,000*$40 = $1,000,000.